# Conjugate Relation Modeling for Few-Shot Knowledge Graph Completion

### v1

@@ 2.1 Neighborhood Aggregation Encoder
To expand the receptive field of the target node and suppress distant neighbor noise, we propose a neighborhood aggregation encoder that combines a gating mechanism with RGAT. For the neighbor set $\mathcal{N}_{h}$ of target node $h$, each neighbor entity $\mathbf{e}_{i}$ is concatenated with its relation embedding $\mathbf{r}_{i}$ and mapped into a neighborhood vector $\mathbf{m}_{i}=W_{r}[\mathbf{e}_{i};\mathbf{r}_{i}]+b_{r}$, where $W_{r}$ and $b_{r}$ are trainable parameters. The aggregated neighborhood representation $\mathbf{h}_{\text{agg}}$ is then computed through attention:
@@
$\mathbf{h}_{\text{agg}}=\sum_{i\in\mathcal{N}_{h}}\alpha_{i}\cdot\mathbf{m}_{i},$ (2)
@@
$\textbf{h}^{{}^{\prime}}=\textbf{g}\odot\textbf{h}_{\text{agg}}+(1-\textbf{g})\odot\textbf{h},\quad\mathbf{g}=\sigma(\textbf{W}_{g}\mathbf{h}_{\text{agg}}+b_{g}),$ (3)
@@ 2.2.1 Implicit Conditional Diffusion Relation Module
where $p_{t}(\textbf{x}_{t}|\textbf{c})$ is the intermediate state distribution conditioned on c, the score function $\nabla{\mathbf{x}_{t}}\log p_{t}(\cdot)$ is approximated by a neural network $\epsilon_{\theta}(\textbf{x}_{t},t,\textbf{c})$ based on a U-Net architecture, and $x_{0}$ is then processed by the attention pool to yield the uncertainty offset $z$.
@@ 2.2.2 Stable Relation Module
Subsequently, attention-weighted pooling is applied to the hidden states of the BiLSTM to generate the final stable relation representation:
$r_{s}=\sum_{i=1}^{K}\frac{\exp(\tanh(\textbf{W}_{s}r_{i}+b_{s}))}{\sum_{j=1}^{K}\exp(\tanh(\textbf{W}_{s}r_{j}+b_{s}))}\,r_{i},$ (11)
@@ 3.1 Experimental Setting
Model NELL-One FB15K-One Wiki-one MRR Hit @ 10 Hit @ 5 Hit @ 1 MRR Hit @ 10 Hit @ 5 Hit @ 1 MRR Hit @ 10 Hit @ 5 Hit @ 1 TransE(2013) 0.168 0.345 0.186 0.082 0.307 0.537 0.419 0.198 0.052 0.090 0.057 0.042 DistMult(2015) 0.214 0.319 0.246 0.140 0.237 0.378 0.287 0.164 0.077 0.134 0.078 0.035 ComplEx(2016) 0.239 0.364 0.253 0.176 0.238 0.370 0.281 0.169 0.070 0.124 0.063 0.030 GMatching(2018) 0.176 0.294 0.233 0.113 0.304 0.456 0.410 0.221 0.263 0.387 0.337 0.197 MetaR(2019) 0.261 0.437 0.350 0.168 0.403 0.647 0.551 0.279 0.221 0.302 0.264 0.178 GANA(2021) 0.344 0.517 0.437 0.246 0.458 0.656 0.575 0.349 0.351 0.446 0.407 0.299 RANA(2023) 0.361 0.573 0.475 0.253 - - - - 0.379 0.480 0.437 0.329 NP-FKGC(2023) 0.460 0.494 0.471 0.437 0.538 0.671 0.593 0.476 0.503 0.668 0.599 0.423 UFKGC(2024) 0.417 0.588 0.511 0.324 - - - - 0.431 0.540 0.491 0.375 PARE(2025) 0.451 0.503 0.458 0.365 0.502 0.664 0.589 0.428 0.478 0.553 0.583 0.429 StarRing(2025) 0.421 0.568 0.512 0.340 0.396 0.598 0.518 0.300 0.406 0.551 0.488 0.330 CR-FKGC 0.534 0.568 0.540 0.516 0.595 0.752 0.696 0.508 0.561 0.725 0.638 0.484
@@ 3.2 Main Results and Analysis
Table 1 shows the performance comparison between CR-FKGC and baseline methods. For fair comparison, all results are reported from the original papers. Specifically, CR-FKGC outperforms state-of-the-art methods, achieving gains of 7.4%, 5.7%, and 5.8% in MRR and 7.9%, 3.2%, and 6.1% in Hits@1, while also attaining superior performance on Hits@5 and Hits@10. This demonstrates that the proposed model not only outperforms the overall ranking performance but also has stronger discriminative power in tail entity recognition accuracy. This performance improvement is primarily due to the synergistic effect of the conjugate relation learner and the neighborhood aggregation encoder: the former effectively alleviates the semantic ambiguity of limited samples by simultaneously modeling relational stable semantics and uncertainty offsets; the latter expands the entity receptive field and suppresses redundant noise, improving semantic integrity. In addition, MaConE maps triples into a manifold space, further enhancing the ability to model complex relations. Notably, CR-FKGC achieves more significant improvements than uncertainty modeling methods such as NP-FKGC and UFKGC. This is due to both the uncertainty modeling advantage of the diffusion model in the denoising process and the implicit conditional embedding generated by the neural process, which makes the model more adaptable under limited samples.
@@
Variants NELL-One FB15K-One MRR Hit@10 Hit@1 MRR Hit@10 Hit@1 w/o Gate 0.512 0.564 0.484 0.559 0.743 0.466 w/o Condition 0.513 0.550 0.485 0.548 0.737 0.468 w/o ICDR 0.504 0.547 0.475 0.536 0.679 0.456 w/o SR 0.470 0.498 0.453 0.510 0.674 0.436 w/o MaConE 0.528 0.554 0.510 0.586 0.732 0.494 CR-FKGC 0.534 0.568 0.516 0.595 0.752 0.508

### v2

@@ 2.1 Neighborhood Aggregation Encoder
To expand the receptive field of the target node and suppress distant neighbor noise, we propose a neighborhood aggregation encoder that combines a gating mechanism with RGAT. For the neighbor set $\mathcal{N}_{h}$ of target node $h$, each neighbor entity $\mathbf{e}_{i}$ is concatenated with its relation embedding $\mathbf{r}_{i}$ and mapped into a neighborhood vector $\mathbf{m}_{i}=W_{r}[\mathbf{e}_{i};\mathbf{r}_{i}]+b_{r}$, where $W_{r}$ and $b_{r}$ are trainable parameters. The aggregated neighborhood representation $\mathbf{h}_{\text{a}}$ is then computed through attention:
@@
$\mathbf{h}_{\text{a}}=\sum_{i\in\mathcal{N}_{h}}\alpha_{i}\cdot\mathbf{m}_{i},$ (2)
@@
$\textbf{h}^{{}^{\prime}}=\text{ReLU}\big(\textbf{g}\odot\textbf{h}_{\text{a}}+(1-\textbf{g}\big)\odot\textbf{h}),\quad\mathbf{g}=\sigma(\textbf{W}_{g}\mathbf{h}_{\text{a}}+b_{g}),$ (3)
@@ 2.2.1 Implicit Conditional Diffusion Relation Module
where $p_{t}(\textbf{x}_{t}|\textbf{c})$ is the intermediate state distribution conditioned on c, the score function $\nabla{\mathbf{x}_{t}}\log p_{t}(\cdot)$ is approximated by a neural network $\epsilon_{\theta}(\textbf{x}_{t},t,\textbf{c})$ based on a U-Net architecture, and $x_{0}$ is then processed by the attention pooling to yield the uncertainty offset $z$.
@@ 2.2.2 Stable Relation Module
Subsequently, we employ a attention pooling and a linear layer to obtain the stable relation embedding:
$\textbf{r}_{s}=\text{FC}(\hat{\textbf{r}}),\quad\hat{\textbf{r}}=\sum_{i=1}^{K}\frac{\exp(\tanh(\textbf{W}_{s}r_{i}+b_{s}))}{\sum_{j=1}^{K}\exp(\tanh(\textbf{W}_{s}r_{j}+b_{s}))}\,r_{i},$ (11)
@@ 3.1 Experimental Setting
Model NELL-One FB15K237-One Wiki-one MRR Hit @ 10 Hit @ 5 Hit @ 1 MRR Hit @ 10 Hit @ 5 Hit @ 1 MRR Hit @ 10 Hit @ 5 Hit @ 1 TransE(2013) 0.168 0.345 0.186 0.082 0.307 0.537 0.419 0.198 0.052 0.090 0.057 0.042 DistMult(2015) 0.214 0.319 0.246 0.140 0.237 0.378 0.287 0.164 0.077 0.134 0.078 0.035 ComplEx(2016) 0.239 0.364 0.253 0.176 0.238 0.370 0.281 0.169 0.070 0.124 0.063 0.030 GMatching(2018) 0.176 0.294 0.233 0.113 0.304 0.456 0.410 0.221 0.263 0.387 0.337 0.197 MetaR(2019) 0.261 0.437 0.350 0.168 0.403 0.647 0.551 0.279 0.221 0.302 0.264 0.178 GANA(2021) 0.344 0.517 0.437 0.246 0.458 0.656 0.575 0.349 0.351 0.446 0.407 0.299 RANA(2023) 0.361 0.573 0.475 0.253 - - - - 0.379 0.480 0.437 0.329 NP-FKGC(2023) 0.460 0.494 0.471 0.437 0.538 0.671 0.593 0.476 0.503 0.668 0.599 0.423 UFKGC(2024) 0.417 0.588 0.511 0.324 - - - - 0.431 0.540 0.491 0.375 PARE(2025) 0.451 0.503 0.458 0.365 0.502 0.664 0.589 0.428 0.478 0.553 0.583 0.429 StarRing(2025) 0.421 0.568 0.512 0.340 0.396 0.598 0.518 0.300 0.406 0.551 0.488 0.330 CR-FKGC 0.534 0.568 0.540 0.516 0.595 0.752 0.696 0.508 0.561 0.725 0.638 0.484
@@ 3.2 Main Results and Analysis
Table 1 shows the performance comparison between CR-FKGC and baseline methods. For fair comparison, all results are reported from the original papers. Specifically, CR-FKGC outperforms state-of-the-art methods across all three datasets, achieving gains of 7.4%, 5.7%, and 5.8% in MRR and 7.9%, 3.2%, and 6.1% in Hits@1, while also attaining superior performance on Hits@5 and Hits@10. This demonstrates that the proposed model not only outperforms the overall ranking performance but also has stronger discriminative power in tail entity recognition accuracy. This performance improvement is primarily due to the synergistic effect of the conjugate relation learner and the neighborhood aggregation encoder: the former effectively alleviates the semantic ambiguity of limited samples by simultaneously modeling relational stable semantics and uncertainty offsets; the latter expands the entity receptive field and suppresses redundant noise, improving semantic integrity. In addition, MaConE maps triples into a manifold space, further enhancing the ability to model complex relations. Notably, CR-FKGC achieves more significant improvements than uncertainty modeling methods such as NP-FKGC and UFKGC. This is due to both the uncertainty modeling advantage of the diffusion model in the denoising process and the implicit conditional embedding generated by the neural process, which makes the model more adaptable under limited samples.
@@
Variants NELL-One FB15K237-One MRR Hit@10 Hit@1 MRR Hit@10 Hit@1 w/o Gate 0.512 0.564 0.484 0.559 0.743 0.466 w/o Condition 0.513 0.550 0.485 0.548 0.737 0.468 w/o ICDR 0.504 0.547 0.475 0.536 0.679 0.456 w/o SR 0.470 0.498 0.453 0.510 0.674 0.436 w/o MaConE 0.528 0.554 0.510 0.586 0.732 0.494 CR-FKGC 0.534 0.568 0.516 0.595 0.752 0.508